1. The blood groups of 200 people are distributed as follows: 40 have type A blood, 75 have type B blood, 60 have type O blood, and 25 have type AB blood. If a person from any of the group is selected at random, what is the probability that this person has an O blood type? How about the AB blood type?
1
Expert's answer
2022-04-24T18:11:08-0400

P(O)=60/200=0.3

P(AB)=25/200